# Omni-Thinker: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards

### v2

@@ 5.1 Training Datasets
We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation. For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning. We further subsample 12,000 examples to fit our compute budget. For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed. Each entry includes a code prompt and JSON-formatted unit tests for automatic validation. For General QA, inspired by SuperGPQA (M-A-P, 2025), we subsample 5,500 queries from the benchmark dataset, proportionally by question category. Each sample comprises a factual question paired with a plain-text answer. We then generate 15 additional confusion options while making sure the uniqueness of correctness by prompting an LLM. The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on single-turn completions. Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
5.2 Baselines
We adopt Qwen2.5-7b-Instruct as the base model for all our experiments (Yang et al., 2024). Its robust instruction-following ability makes it a suitable candidate for subsequent reinforcement learning training on both reasoning tasks and more general open-domain QA, as it relies on the model’s capacity to comprehend and follow the given prompts effectively.
Supervised Fine-Tuning (SFT):
@@ 5.2 Baselines
In order to have a meaningful comparison with GRPO, we adopt a similar self-sampled data curation and fine-tuning approach with Rejection sampling Fine-Tuning (Yuan et al., 2023). We first prompt the base model to generate 128 chain-of-thought responses for our training dataset to ensure we end up with at least one correct response for most queries, then filter them based on the same accuracy reward signals used in GRPO training. We then perform supervised fine-tuning on Qwen2.5-7b-Instruct using these self-distilled responses. This provides a strong on-policy learning baseline that incorporates explicit reasoning steps through self-distillation from the base model.
Model Merging:
We employ TIES-Merging  (Yadav et al., 2023b) as our model‑merging baseline. It’s a simple yet effective method designed specifically for the multi-task setting that takes into consideration the interference between parameters from models trained on individual tasks during the merging process. It has demonstrated superior performance in multi-task learning compared to linear and task arithmetic approaches (Yadav et al., 2023a). To begin with, we conduct single-task GRPO training using individual task datasets and collect the model weights of the best checkpoints with the help of a validation set for each training run. We then merge the four single-task models using a scaling value $\lambda=1$.
5.3 Evaluation
We assess performance in each domain using dedicated held-out benchmarks aligned with the task’s unique evaluation criteria. The detailed evaluation set is presented as follows
Math Reasoning:
we compute accuracy over seven datasets: AIME24 (MAA, 2024), AMC23 (MAA, 2023), Gaokao2023EN (Liao et al., 2024), MATH-500 (Hendrycks et al., 2021), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024).
Code Generation:
we measure coding ability via pass@1 on BigCodeBench (Complete-Full) (Zhuo et al., 2024) and LiveCodeBench (24Oct-25Jan) (Jain et al., 2024).
General QA:
we report exact-match accuracy using the MMLU-Pro benchmark (Wang et al., 2024).
Creative Writing:
we evaluate the creative writing task using the role-play and creative writing subcategory of MT‑Bench (Zheng et al., 2023), reporting win rate against a GPT-4 (pre-gen dated June 16, 2023) model.
@@ 6 Results and Discussion
Figure 2: Performance gains across four task domains, comparing our Multi-Task (MT) framework (Joint and Curriculum variants) against baselines including Supervised Fine-Tuning (SFT) and Model Merging. Curriculum MT achieves the strongest results, particularly in open-ended tasks, showing that controlling how multi-task learning is structured is crucial for effective generalization.
@@ 6.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
We evaluate Omni-Thinker across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization. Figure 2 shows that Curriculum-Guided Multi-Task GRPO consistently yields the best results. Table 2 further details how these gains vary by benchmarks.
In Math, Curriculum MT achieves the highest average performance at 59.6%, with the clearest gains on more complex reasoning tasks such as MinervaMath and OlympiadBench. These benchmarks benefit from strong rule-based reward signals and backward-transfer-guided task ordering. In contrast, datasets like AMC23 show minimal change because their relatively high baseline scores likely reflect smaller question sets and potential pretraining overlap rather than robust multi-step problem-solving.
In General QA, Curriculum MT again performs best (52.2%), followed by Model Merging (49.8%) and Joint Multi-Task GRPO (48.8%). These improvements are driven by our Short-Form Open-Ended Supervision strategy: instead of generating responses in a fully open-ended and unconstrained fashion, the model is trained to produce complete answer strings given a diverse set of candidate responses, enabling the effective application of verifiable reward through simple string matching when training general-domain tasks.
For Code Generation, Curriculum MT achieves 35.4%, slightly ahead of Model Merging. Notably, we only evaluate on the subset of LiveCodeBench(24Oct-25Jan) problems released after Qwen2.5’s data cutoff, which ensures that these are unseen test items. This setup highlights Curriculum MT’s significant generalization gains on novel problems, explaining the larger improvements on LiveCodeBench relative to static benchmarks like BigCodeBench, where data overlap is more likely.
@@
These results support our central hypothesis: The Omni-Thinker Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
Figure 3: Average backward transfer (BWT) per target task, showing how training single-task GRPO on all other tasks affects it on average. A negative BWT indicates forgetting; a positive value indicates positive transfer.
6.2 Training Order Matters: From Least to Most Forgettable
Our analysis highlights that task ordering is critical for robust multi-task reinforcement learning. Figure 4 summarizes the average backward transfer (BWT) each task receives when training single-task GRPO models on other tasks. Specifically, the x-axis of Figure 4 represents the target task. A negative average BWT value for a target task indicates that it tends to suffer forgetting or experience a decrease in performance when other tasks are trained before it; a positive value suggests that the target task tends to benefit or experience positive transfer. This high-level view reveals that Creative Writing is the most vulnerable task, as it experiences the most negative average BWT received, while Coding is the most resilient, showing a positive average BWT received.
For example, Coding benefits on average after training on another task (average BWT received $+1.06\%$, indicating a positive transfer). In contrast, Creative Writing is the most vulnerable to forgetting: it has the highest magnitude of negative average BWT received ($-4.57\%$), meaning it easily loses performance when other tasks are trained first.
Figure 5 shows the full BWT matrix, where x-axis consists of target tasks and y-axis represent source tasks that are being trained on. Each cell $(row_{i},col_{j})$ in the matrix represents the BWT from source task $j$ to target task $i$. A negative value indicates that training on source task $j$ causes forgetting in target task $i$, while a positive value suggests positive transfer. We observe that structured tasks like Coding and Math are both less forgetting conducive as sources and tend to receive more positive transfers as targets from training on other tasks.
Guided by this analysis, our curriculum trains tasks in a least-to-most forgettable order: Coding $\rightarrow$ Math $\rightarrow$ General QA $\rightarrow$ Creative Writing. This progression grounds the model in stable, verifiable tasks before introducing those more prone to interference and forgetting. As confirmed by Table 1, reversing this order significantly harms performance.
Overall, our findings suggest that BWT patterns can offer a practical, empirically informed guide for task sequencing, helping reduce negative transfer and improve generalization in multi-task RL for LLMs. Further work is needed to test this approach across a broader range of tasks and domains, including those that require logical reasoning over graph-structured data  (Zhou et al., 2024) and knowledge-base retrieval  (Dehghan et al., 2024), as well as more diverse open-ended domains such as mixed-initiative collaborative storytelling and co-creativity  (Kreminski et al., 2024).
6.3 Output Format Matters: Full-Text Answers Enhance Generalization
We examine how output format impacts generalization by comparing models trained to generate full-text answers versus selecting letter choices in multiple-choice QA (MCQ). Using GRPO, we train two single-task policies on the training set, one prompted to produce full-text final answers at the end of its chain-of-thought completions, and the other to output only letter choices (e.g., “A”, “B”, “C”). As shown in Figure 4, the model trained to output full-text answer achieves significantly better generalization when evaluated with free-form QA prompts on MMLU Pro (51% vs. 41%). While the letter-choice model slightly outperforms when evaluated strictly on MCQ prompts, the full-text model remains competitive across both prompt formats.
@@ A.1 Reward Estimation
Omni-Thinker employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework. We define task-specific reward functions as $r_{k}(o,\phi_{k})$, where $o$ denotes the model output and $\phi_{k}$ represents auxiliary task-specific inputs required for evaluation—such as reference answers, unit tests, or candidate responses for comparison. Each reward function captures domain-relevant correctness criteria, assessing whether $o$ satisfies symbolic constraints, passes execution tests, or is preferred over alternatives under subjective evaluation. While some rewards (e.g., math and code) are strictly deterministic, others, such as LLM-as-a-Judge comparisons, are inherently stochastic but executed at low decoding temperature to ensure stable and consistent supervision. All reward functions are designed to be domain-aware and automatable, supporting scalable reinforcement learning across both structured and generative tasks.
@@
  • Math:
    $r_{\texttt{math}}(o,a)=\mathbbm{1}\left\{\texttt{verify}_{\texttt{math}}(o_{\texttt{ans}},a)=\texttt{true}\right\}\,,$
    where $a$ denotes the ground-truth answer, and $o_{\texttt{ans}}$ is the predicted answer extracted from the <answer> tags in the model output $o$. Symbolic equivalence between $o_{\texttt{ans}}$ and $a$ is verified using a deterministic parser.
  • Code Generation:
[…]
    where the generated code $o_{\texttt{ans}}$ is executed in a sandboxed environment and evaluated against the unit tests defined by $x$; $\models$ indicates logical satisfaction.
  • General QA:
    $r_{\texttt{qa}}(o,a)=\mathbbm{1}\left\{o_{\texttt{ans}}=a\right\}$
@@
  • Creative Writing:
    $r_{\text{writing}}(o,o_{ref})=\begin{cases}1.0&\text{if }o\succ o_{ref}\\ 0.5&\text{if }o\sim o_{ref}\\ 0.0&\text{if }o\prec o_{ref}\end{cases}$
    where $o\succ o_{ref}$ indicates that the model output $o$ is preferred over the reference $o_{ref}$, $o\prec o_{ref}$ means the reference is preferred, and $o\sim o_{ref}$ denotes a tie. Preferences are judged via pairwise comparison by a fixed LLM-as-a-Judge model.
@@
$r_{\texttt{format}}(o)=\mathbbm{1}\left\{\texttt{tags\_valid}(o)\right\}$
$r_{\texttt{tags}}(o)=\frac{1}{4}\cdot\left|\texttt{tags\_present}(o)\right|$
@@
$\mathcal{R}_{k}(o,\phi_{k})=\sum_{r\in\mathcal{F}_{k}}w_{r}\cdot r(o,\phi_{k}),$
where $w_{r}\in[0,1]$ denotes the task-specific weight for reward component $r$. If a component reward is undefined, e.g., due to malformed or unparsable output, it is omitted from the sum. Samples with no valid reward components are excluded from policy updates.

### v3

@@ 4 Experimental Setup
We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation. For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning. We further subsample 12,000 examples to fit our compute budget. For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed. Each entry includes a code prompt and JSON-formatted unit tests for automatic validation. For General QA, we subsample 5,500 queries from from SuperGPQA (Du et al., 2025) dataset, proportionally by question category. Each sample comprises a factual question paired with a plain-text answer. We then generate 15 additional confusion options while making sure the uniqueness of correctness by prompting an LLM. The Creative Writing. domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on single-turn completions. Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
Evaluation.
We assess performance in each domain using dedicated benchmarks aligned with the task’s evaluation criteria. Math: accuracy on AIME24 (MAA, 2024), AMC23 (MAA, 2023), Gaokao2023EN (Liao et al., 2024), MATH-500 (Hendrycks et al., 2021), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). Coding: pass@1 on BigCodeBench (Complete-Full) (Zhuo et al., 2024) and LiveCodeBench (24Oct–25Jan) (Jain et al., 2024). General QA: exact-match accuracy on MMLU-Pro (Wang et al., 2024). Creative Writing: win rate on the role-play and creative writing subsets of MT-Bench (Zheng et al., 2023), against GPT-4 (pre-gen, June 16, 2023).
Baselines.
@@
We use Qwen2.5-7B-Instruct (Yang et al., 2024) as the base model for all experiments, owing to its strong instruction-following capability, which makes it well-suited for reinforcement learning on both structured reasoning and open-domain QA tasks. Supervised Fine-Tuning (SFT): In order to have a meaningful comparison with GRPO, we adopt a similar self-sampled data curation and fine-tuning approach with Rejection sampling Fine-Tuning (Yuan et al., 2023). We first prompt the base model to generate 128 chain-of-thought responses for our training dataset to ensure we end up with at least one correct response for most queries, then filter them based on the same accuracy reward signals used in GRPO training. We then perform sft on base model using these self-distilled responses. This provides a strong on-policy learning baseline that incorporates explicit reasoning steps through self-distillation from the base model. Model Merging: We employ TIES-Merging  (Wu et al., 2025) as our model‑merging baseline. It is a simple yet effective method designed specifically for the multi-task setting that takes into consideration the interference between parameters from models trained on individual tasks during the merging process. It has demonstrated superior performance in multi-task learning compared to linear and task arithmetic approaches (Yadav et al., 2023). To begin with, we conduct single-task GRPO training using individual task datasets and collect the model weights of the best checkpoints with the help of a validation set for each training run. We then merge the four single-task models using a scaling value $\lambda=1$.
@@ 5.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
We evaluate Omni-Thinker across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization. BWT matrix is computed following equation 6, then Algorithm 1 is used to predict the accuracy of the model after curriculum learning, Appendix A.2.2 for details. The predicted best curriculum using Algorithm 2 is Code $\rightarrow$ Math $\rightarrow$ QA $\rightarrow$ Writing while the worst is Writing $\rightarrow$ QA $\rightarrow$ Math $\rightarrow$ Coding.
Figure 1 shows that Curriculum Multi-Task Learning with GRPO consistently yields the best results. Table 1 further details how these gains vary by benchmarks.
In Math, Curriculum Learning (CL) achieves the highest average performance at 59.6%, with the clearest gains on more complex reasoning tasks such as MinervaMath and OlympiadBench. These benchmarks benefit from strong rule-based reward signals and backward-transfer-guided task ordering. In contrast, datasets like AMC23 show minimal change because their relatively high baseline scores likely reflect smaller question sets and potential pretraining overlap rather than robust multi-step problem-solving.
In General QA, CL again performs best (52.2%), followed by Model Merging (49.8%) and Mixed Training GRPO (48.8%). These improvements are driven by our Short-Form Open-Ended Supervision strategy: instead of generating responses in a fully open-ended and unconstrained fashion, the model is trained to produce complete answer strings given a diverse set of candidate responses, enabling the effective application of verifiable reward through simple string matching when training general-domain tasks.
For Code Generation, CL achieves 35.4%, slightly ahead of Model Merging. Notably, we only evaluate on the subset of LiveCodeBench(24Oct-25Jan) problems released after Qwen2.5’s data cutoff, which ensures that these are unseen test items. This setup highlights CL’s significant generalization gains on novel problems, explaining the larger improvements on LiveCodeBench relative to static benchmarks like BigCodeBench, where data overlap is more likely.
@@
These results support our central hypothesis: The Omni-Thinker Framework, BWT-guided Curriculum Learning with hybrid rewards, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
Table 1: Performance across benchmarks. ST = Single-Task RL (e.g., ST Math = RL trained only on math). $M$ = Model Merging. JT = Joint Training. CL = Curriculum Learning. Bolded values mark the best per row; underscored values mark the second best. Domains include Math (7 sets), MMLU-Pro (9 categories), Coding (2 sets), and Creative Writing (MT-Bench).
[…]
Math
AIME24 18.0 13.3 14.7 14.0 15.3 16.7 10.0 11.3 15.3
AMC23 57.5 57.5 60.0 62.5 61.0 62.0 56.0 51.0 70.0
Gaokao2023en 73.0 74.3 76.1 74.0 75.6 74.3 74.8 76.6 77.1
MATH500 78.2 78.8 80.4 75.4 79.2 76.8 79.8 77.6 81.0
MinervaMath 64.3 64.0 66.5 63.2 61.8 65.1 66.2 68.4 71.7
OlympiadBench 42.1 43.0 43.7 41.3 43.0 43.0 41.8 43.6 47.4
Average 55.5 55.1 56.9 55.1 56.0 56.3 54.8 54.8 60.4
General QA
Biology 57.6 56.8 52.3 67.4 59.0 66.3 65.6 67.2 68.8
Business 33.5 39.0 25.6 58.7 33.0 48.2 59.8 49.8 47.5
Chemistry 35.8 31.8 27.3 47.7 38.3 44.1 42.5 42.1 50.7
CS 53.7 48.1 50.2 55.1 52.0 53.7 53.9 58.8 59.3
Economics 42.7 49.2 38.7 62.9 44.9 59.6 62.0 56.8 62.1
Engineering 28.3 31.3 20.4 37.5 26.6 37.8 38.1 35.8 37.1
Health 46.7 46.2 45.2 51.0 47.2 45.7 52.7 50.7 57.1
History 37.3 33.3 34.7 47.2 38.6 33.9 47.3 43.3 45.7
Law 23.2 24.0 20.6 27.9 23.3 26.8 26.6 27.5 29.7
Math 55.4 52.6 50.4 59.3 56.3 57.4 58.3 59.2 61.2
Other 44.3 40.0 39.7 51.0 43.9 46.4 51.8 49.9 53.3
Philosophy 36.9 34.3 33.3 43.9 35.5 38.2 41.5 42.1 42.9
Physics 41.1 37.4 30.8 53.7 41.6 49.8 46.7 48.0 55.6
Psychology 50.9 51.5 45.4 60.2 51.8 59.0 59.4 59.3 61.8
Average 41.5 40.1 37.9 51.3 42.0 47.8 49.8 48.8 52.2
Coding
BigCodeBench 46.5 50.4 46.7 47.1 46.8 44.5 48.1 47.2 49.5
LiveCodeBench 12.7 21.8 13.1 13.8 13.3 14.2 20.8 17.0 21.3
Average 29.6 36.1 29.9 30.5 30.1 29.3 34.4 32.1 35.4
Creative Writing
MT-Bench (Writing) 74.2 71.6 74.2 63.0 78.3 44.2 67.5 83.0 84.2
5.2 Entropy Dynamics: Discussion
Comparison between accuracies predictions to the actual test evaluation results for various curricula is depicted on Table 3. Predicted accuracies using test set backward transfers are surprisingly precise considering our simplifying assumptions, especially for the top curriculum. We now discuss an identified cause of discrepancies.
We define the token-wise entropy of a policy $\pi_{\theta}$ on task $T_{k}$ ti measures the average per-token uncertainty of the policy across task samples.
$\mathcal{E}(\theta,T_{k}):=-\mathbb{E}_{q\sim\mathcal{D}_{k},o\sim\pi_{\theta}^{\otimes}(\cdot\mid q)}\frac{1}{|o|}\sum_{t=1}^{|o|}\sum_{v\in\mathcal{V}}\pi_{\theta}(v\mid q,o_{<t})\log\pi_{\theta}(v\mid q,o_{<t}).$ (11)
Table 2: Test performance (%) of single-task RL fine-tuning on General QA and Creative Writing respectively under different generation temperature (T) in training.
Eval Task General QA Creative Writing
T=1.0 T=1.2 T=1.0 T=0.1
Math 55.1 54.8 55.6 57.8
Coding 30.4 26.6 30.1 27.4
General QA 51.4 13.9 42.0 44.9
Creative Writing 63.0 66.7 78.3 82.5
Entropy has been shown to drop during GRPO fine-tuning on reasoning, coding and more generally verified rewards (Rastogi et al., 2025; Cui et al., 2025; Yu et al., 2025) with a correlation to higher accuracy until reaching a breaking point. Long training requires extra care towards entropy scaling either via regularization (Shen, 2025) or dynamic temperature scaling. Our multi-task setting differs in two key aspects compared to the above references.
First, we are using hybrid rewards with both verified and generative components. The Creative Writing task is generative and is expected to increase entropy. Indeed, Wang et al. (2025) account for so-called forking tokens corresponding to structural choices of the output capturing most of the entropy. Reasoning tasks tend to require highly causal token sequences hence few forking tokens (low entropy) while generative tasks may allow more logical cuts at inference (higher entropy).
Second, we train on multiple domains compared to mostly single-domain analysis in the references above. It is unclear a priori whether entropy decrease propagates from task to task. Agarwal et al. (2025) show that fine-tuning to reduce entropy suffices to improve performance on multiple domains. We hypothesize that models implicitly learn to emulate lower or higher temperatures as a mechanism to regulate entropy. In practice, the policy often produces logically flawed but rarely syntactically meaningless outputs, suggesting that its support lies within a constrained domain $V(q,o_{<t})\subset\mathcal{V}$: $\sum_{v\in V(q,o_{<t})}\pi_{\theta}(v\mid q,o_{<t})=1$. Scaling the final layer weights by a factor $\lambda<1$ preserves this domain while increasing entropy, effectively raising the model’s base temperature. Such changes propagate across all tasks, not only the one being fine-tuned. Thus, even when two domains are sufficiently distinct for knowledge transfer to fail, the entropy dynamics may still be measurable across tasks.
Table 3: Comparison of empirical and predicted test accuracies (%). Each task column reports Test vs. Predicted accuracy for a given curriculum order. Standard deviations are rounded up.
Curriculum Math Coding QA Writing
Test Pred Test Pred Test Pred Test Pred
$\mathcal{C}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{W}$ $60.4\!\pm\!0.3$ $57.3$ $\pm\!0.3 35.435.4$ $52.2\!\pm\!0.1$ $51.9$ $84\!\pm\!2$ $78.4$
$\mathcal{Q}\,\mathcal{M}\,\mathcal{W}\,\mathcal{C}$ $9.3\!\pm\!0.3 55$.655.6 $31.6\!\pm\!0.3$ $36.7$ $39.0\!\pm\!0.1$ $38.8$ $79\!\pm\!2$ $78.4$
$\mathcal{Q}\,\mathcal{W}\,\mathcal{C}\,\mathcal{M}$ $60.4\!\pm\!0.3$ $57.6$ $1.9\!\pm\!0.3 33$.933.9 $36.3\!\pm\!0.1$ $38.8$ $82\!\pm\!2$ $73.1$
$\mathcal{W}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{C}$ $6.6\!\pm\!0.3 55$.555.5 $32.7\!\pm\!0.3$ $36.1$ $22.6\!\pm\!0.1$ $38.4$ $75\!\pm\!2$ $62.1$
5.3 Entropy Dynamics: Empirical Support
Figure 2: Validation Set Entropy Change Matrix.
The intuitions laid out in the previous section are empirically supported by two experiments.
We define the entropy change matrix as
$H_{ij}:=\frac{\mathcal{E}(\theta(\theta_{0},T_{i}),T_{j})}{\mathcal{E}(\theta_{0},T_{j})}-1$ (12)
and compute it, see Figure 2. We observe that Math and Coding decrease entropy for all tasks (as previously observed for Verified Rewards) while Creative Writing increases entropy. Also, entropy change seems to depend primarily on the source task type, secondarily on the target task type, but not on their domain overlap.
We fine-tune the base model on QA and Writing task with different choices of temperature to emulate the effect of entropy modifications due to fine-tuning on entropy-increasing or entropy-decreasing tasks. On the one hand, Writing is thought to benefits from temperature lowering coming from other tasks, we thus train the base model on writing with a lower temperature and expect the model to close the gap compared to the best Curriculum-trained model when evaluated with zero temperature. On the other hand, QA is thought to performs worse than expected in the worst curriculum due to the increased entropy coming from the trainig on Writing. We fine-tune the base model on QA with a higher temperature, then evaluate with zero temperature, and expect QA performance to drop toward the low performance of the worse curriculum learning. The results on table 2 shows that indeed the case.
6 Related Work & limitations
Large Language Models and Multi-Task Learning
Large Language Models and Reinforcement Learning
@@ 7 Conclusion
Limitations:
Further work is needed to test this approach across a broader range of tasks and domains, including those that require logical reasoning over graph-structured data  (Zhou et al., 2024) and knowledge-base retrieval  (Dehghan et al., 2024), as well as more diverse open-ended domains such as mixed-initiative collaborative storytelling and co-creativity  (Kreminski et al., 2024). Our discussion on entropy is consistent with the existing literature however we lack 1) experiments backing our intuition that forking tokens are clearly identifiable and more frequent in generative tasks is lacking 2) a direct measurement of temperature scaling due to fine-tuning. Furthermore, our discussion on entropy is mostly qualitative. A quantitative study of entropy effects could allow to avoid failure cases of our accuracy predictions hence of our task scheduler. Finally, we fully leverage BWT only for curriculum learning. A comprehensive framework combining BWT, entropy and gradient norm in a quantitative manner in the continuous limit would allow for dynamical joint training with mixed batches.
@@ A.1 Reward Estimation
Omni-Thinker employs a hybrid reward system combining rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework. We define task-specific reward functions as $R_{k}(q,o)$, where $o$ denotes the model output and $q$ is the prompt provided to the model. Each reward function captures domain-relevant correctness criteria, assessing whether $o$ satisfies symbolic constraints, passes execution tests, or is preferred over alternatives under subjective evaluation. While some rewards (e.g., math and code) are strictly deterministic, others, such as LLM-as-a-Judge comparisons, are inherently stochastic but executed at low decoding temperature to ensure stable and consistent supervision. All reward functions are designed to be domain-aware and automatable, supporting scalable reinforcement learning across both structured and generative tasks.
@@
  • Math: Elements of the math dataset are couples $(q,a_{q})$ where $q$ is a prompt and $a_{q}$ is a token sequence stating the answer. We implement a $\texttt{verify}_{\texttt{math}}(o,a)$ function that combines regular expression and symbolic parser to checks that $a$ (or an equivalent acceptable answer) is in $o$ within tags <answer>. More formally
    $r_{\texttt{math}}(q,o)=\mathbbm{1}\left\{\texttt{verify}_{\texttt{math}}(o,a_{q})=\texttt{true}\right\}\,.$
  • Code Generation: Each element of the code dataset is a tuple $(q,\texttt{unittest}_{q},\texttt{test\_case}_{q})$ where $q$ is a prompt, $\texttt{unittest}_{q}$ is a unit test function and $\texttt{test\_case}_{q}$ is a set of test cases. Given an output $o$ for prompt $q$, the generated code $o_{\texttt{ans}}$ is extracted from the output $o$ using regular expressions, the unit test $\texttt{unittest}_{q}(o_{\texttt{ans}},x)$ is executed in a sandboxed environment for every test case $x\in\texttt{test\_case}_{q}$. More formally
    $r_{\texttt{code}}(q,o)=\prod_{x\in\texttt{test\_case}_{q}}\mathbbm{1}\left\{\texttt{unittest}_{q}(o_{\texttt{ans}},x)\right\}$
  • General QA: Each element of the dataset for General QA is a couple $(q,a_{q})$ of prompt and answer. The reward is defined by extracting the answer $o_{\texttt{ans}}$ from the output $o$ using regular expressions and testing it against the ground truth $a$. More formally:
    $r_{\texttt{qa}}(q,o)=\mathbbm{1}\left\{o_{\texttt{ans}}=a_{q}\right\}$
@@
  • Creative Writing: Each element of the dataset for Creative writing is a couple $(q,o_{\texttt{ref},q})$. Given an output $o$ on prompt $q$, the reward is computed by calling a fixed LLM-as-a-Judge model prompted to do a pairwise comparison between $o$ and $o_{\texttt{ref},q}$. More formally:
    $r_{\text{writing}}(q,o)=\begin{cases}1.0&\text{if }o\succ_{q}o_{\texttt{ref},q}\\ 0.5&\text{if }o\sim_{q}o_{\texttt{ref},q}\\ 0.0&\text{if }o\prec_{q}o_{\texttt{ref},q}\end{cases}$
    where $A\succ_{q}B$ means that the fixed LLM-as-a-Judge model prefered $A$ over $B$ for request $q$, and $A\sim_{q}B$ means it judges the answers as tied.
@@
$r_{\texttt{format}}(q,o)=\mathbbm{1}\left\{\texttt{tags\_valid}(o)\right\}$
$r_{\texttt{tags}}(q,o)=\frac{1}{4}\cdot\left|\texttt{tags\_present}(o)\right|$
@@
$R_{k}(q,o)=\sum_{r\in\mathcal{F}_{k}}w_{r}\cdot r(q,o),$
where $w_{r}\in[0,1]$ denotes the task-specific weight for the reward component $r$. If a component reward is undefined, e.g., due to malformed or unparsable output, it is omitted from the sum. Samples with no valid reward components are excluded from policy updates.